# Chapter 1      Introduction to IoT

### 09 Dec 2023 v1

@@ 1.2.2 Components of IoT
1.2.3 Characteristics of IoT
Characteristics pertain to the distinct qualities or capabilities of a system that set it apart from others. When it comes to IoT, providing a fixed set of characteristics can be challenging due to the evolving nature of IoT devices and their dependency on the specific application domain. Nevertheless, there are several common traits that help differentiate IoT technology from other systems. This section describes the characteristics of IoT system [31, 32]. Fig. 1.1 provides a visual representation of the characteristics of IoT.
Figure 1.1: IoT characteristics
@@ 1.2.3 Characteristics of IoT
    Connectivity
    Connectivity is the most vital requirement of IoT. The main aspect of Internet of Things is a network with millions of devices connected to each other. The connectivity remains constant, allowing anyone from anywhere to connect to the IoT network at any given moment.
@@
    Intelligent decision making capability
    The extraction of knowledge from the data generated is highly significant. Consider a sensor that produces data; however, the true value of that data lies in its proper interpretation. This represents a crucial aspect of IoT, wherein IoT devices possess the capability to transform raw data collected by sensors into meaningful information and make decisions based on it.
@@
    Dynamic and Self Adapting
    IoT devices should have the ability to adapt to changes in context, their surrounding environment, and the existing situation. For instance, within a surveillance system, cameras are capable of switching between day and night modes or adjusting their resolution in response to motion detection, demonstrating their adaptability.
@@
    Unique Identity
    Every IoT device should have a unique identity and unique identifier. IoT device interfaces enable users to inquire about device information, monitor their status, and remotely manage them. Having a distinct identity is essential to empower users to safeguard their devices, whether through password protection or alternative security measures [33].
@@
    Self Configuring
    IoT devices possess the capability to autonomously update their systems in response to the situation, eliminating the need for user intervention. Moreover, they exhibit flexibility in network management, allowing new devices to seamlessly join the network and permitting any device to depart from the network at any time.
@@
    Scalability
    The IoT network is experiencing a continual growth in the number of connected devices, resulting in a substantial and continuous generation of data. Consequently, scalability emerges as the foremost feature for any IoT system.
    Figure 1.2: Advantages of Internet of Things
@@
    Interoperability
    IoT devices rely on standardized protocols and technologies to guarantee seamless communication among themselves and with other systems. Interoperability represents a fundamental block of the IoT, signifying the capacity for various IoT devices and systems to interact and share data, irrespective of the underlying technology or the manufacturer. Hence, IoT devices employ standardized protocols, data formats and technologies to uphold interoperability.
@@
    Heterogeneity
    The devices within an IoT network exhibit heterogeneity, showcasing the network’s capability to accommodate diverse elements.
@@
    Energy Efficiency
    Energy efficiency stands out as a significant characteristic of IoT. Numerous IoT devices are purposefully crafted to minimize energy consumption and create low-power devices. Furthermore, various IoT approaches are tailored to optimize power usage, such as opting for fog/edge computing over cloud computing to decrease bandwidth requirements and reduce power consumption when constructing an IoT system [34].
@@
    Security
    The presence of millions of internet-connected devices and the vast data generation underscore the vulnerability of the IoT network to security threats. Hence, safety and security emerge as pivotal characteristics of IoT. Ensuring safety is of paramount importance to preserve the effectiveness of IoT’s advantages, including efficiency and novel experiences.
@@ 1.3 Advantages of IoT
The amount of benefit IoT offers to people’s life is the main reason IoT is becoming more popular by the day. IoT solutions are developed and invented to make people’s life easier and more convenient. Healthcare, education, and business—there is hardly any field that is not affected by IoT technology. This section elaborates on the most visible benefits of IoT in people’s daily lives [35, 36] and Fig. 1.2 provides a visual overview.
@@
    Efficient Data Collection
    IoT based data collection has been particularly beneficial in sectors such as healthcare and finance [37]. A straightforward illustration of efficient data collection is the integration of IoT in the retail sector. Internet-connected tags can provide data on purchase decisions and sales trends, whether weekly or monthly. This enhanced data gathering can improve inventory management and reveal valuable insights into customer behavior, ultimately contributing to the prosperity of the business.
@@
    Control and Automation
    IoT has provided its customers a more convenient lifestyle and allows them to control their daily activities with the touch of a button. One simple yet remarkable example is the smart bulb, which can be controlled without even touching the switches; the user can simply turn off or on the light remotely. Not only can the light bulb, coffee maker, or any other electrical device in the home be controlled with the tip of the device, but they can also be controlled using a voice command when they are linked to Google’s or Amazon’s voice assistants.
@@
    Real-time Access to Information
    IoT devices offer immediate access to information, proving invaluable across healthcare, business, and everyday applications. A prime illustration of the advantages of real-time data access is within the healthcare sector. Physicians can access patient data in real-time, enabling continuous health monitoring. This capability becomes particularly crucial in delivering emergency medical assistance swiftly when unexpected health issues arise.
@@
    Improved Efficiency
    IoT systems operate autonomously, a valuable asset in various domains. Reduced human intervention leads to increased efficiency and decreased labor reliance. For instance, a company with a fleet of delivery vehicles can effortlessly track their real-time locations, eliminating the need for manual employee involvement in this task.
@@
    Improved Quality of Life
    The advent of IoT has significantly improved the lives of its users in numerous ways. Real-time health monitoring, including devices like blood pressure monitors and fitness trackers, empowers users to maintain their well-being effectively. Smart homes offer a stress-free and effortless lifestyle. These advantages extend beyond individuals and can benefit entire industries or communities. Smart devices, linked not only to intelligent traffic lights but also road safety monitors and toll gates, can provide drivers with real-time information about road conditions on their route.
@@
    Cost and Time Effective
    IoT minimizes human effort and relies heavily on real-time data transmission, leading to time savings. For example, real-time patient monitoring benefits both patients and doctors by eliminating the need for physical meetings, thus saving time for both parties. IoT aids businesses in streamlining their workflows by offering valuable insights and real-time information, resulting in cost reductions. In addition to businesses, individuals can reduce their everyday expenses through the use of IoT.
@@
    Asset Tracking
    This process involves tracking products within a business or logistics management system. Manual asset tracking is labor-intensive and time-consuming, but it can be streamlined through the application of IoT technologies like barcodes and RFID tags. These technologies allow for remote monitoring of goods and provide stakeholders with information about any faults or problems in real time [38].
@@
    Data Analysis for Research Purpose
    The enormous amount of data collected from the IoT devices has become a blessing for the researchers in various fields like healthcare, education, business etc. The healthcare researchers can use the data collected via biosensors and utilize them in inventing cures and vaccines for disease, the finance industry can utilises the data to understand the trend and improve customer experience, the super shops can analyze the customer behaviour and improve their business and so on.
@@
    Big Data and Predictive Analysis
    “Big data” has been a widely recognized term in the world long before the emergence of IoT. It involves the collection and analysis of massive volumes of data. One of the primary objectives of IoT is to amass data from diverse sources, sending this information back to systems for analysis. Effective analysis of big data can yield valuable insights, spanning from stock market predictions to understanding customer behavior, and thereby enhancing the business landscape.
@@
    Increased Productivity
    IoT utilization in both industry and homes has the potential to boost productivity significantly. For instance, in a smart home, users can streamline various household tasks using voice commands, enabling efficient multitasking. Similarly, in a business setting, analyzing customer behavior can enhance customer satisfaction, ultimately contributing to the prosperity of the enterprise. As an example, 46% of businesses that embraced IoT strategies saw improvements in efficiency, even though only 29% initially anticipated such improvements [39]. In the healthcare sector, doctors can offer more extensive services to their patients if they do not need to make physical visits to each patient individually.
@@
    Safety and Security
    The incorporation of IoT offers users a means of security not only in their homes but also in their businesses, schools, offices, and virtually anywhere. Individuals can remotely monitor their valuable assets such as vehicles and other important items. Parents can even keep track of their children’s whereabouts from their workplaces, providing peace of mind. IoT allows for vehicle tracking and the setup of alert systems in case of unusual incidents. Financial companies and banks can enhance the security of their confidential rooms or vehicles by utilizing IoT [34].
@@
    Enhance Customer Engagement
    The IoT offers several avenues for enhancing customer engagement. It achieves this by leveraging valuable customer data, personalizing experiences, improving convenience, and enabling real-time interactions.
@@
    Efficient Resource Utilization
    IoT facilitates efficient resource utilization through various mechanisms and capabilities. Its ability to collect and monitor real-time information allows organizations to track the status of their resources, such as equipment and machinery, enabling the identification of inefficiencies. Furthermore, in the industrial sector, predictive maintenance powered by IoT can anticipate machinery failures, reducing downtime and optimizing resource allocation for maintenance.
@@
    Improved Technology
    The innovation brought about by IoT leads to the creation of newer and more advanced technologies in the market. For instance, consider the scenario of an air conditioner that was initially controlled manually with a remote. With the advent of IoT, users can now operate it using voice commands or control it remotely. When such innovations hit the market, they spark competition, driving the development of improved solutions based on user feedback. This cycle of innovation in response to IoT advancements contributes to the continuous progression of technology [37].
Along with the generic benefits that apply to practically all fields that are integrating IoT, various domain-specific benefits are also apparent in many sectors. Some of the domains are described below.
Healthcare
The healthcare industry is one of the most benefited by IoT among all the sectors. The innovation of IoT has revolutionized the healthcare industry by various means. Here are some common examples [6]:
@@
    Remote Medical Assistance
    The idea of remote monitoring has brought significant advantages to the healthcare sector. Doctors can now observe their patients’ conditions from their offices without the need for physical visits, just by using the sensors attached to the patients’ bodies.
@@
    Improved Treatment
    Through the utilization of IoT technology, physicians can make well-informed decisions based on evidence and ensure full transparency.
@@
    Faster Disease Diagnosis
    Remote and continuous monitoring allows doctors to identify diseases in their early stages, which can save many lives.
  4. 4.
    Emergency monitoring and reporting
    During remote monitoring, if any unusual or emergency situations arise, such as heart failure or asthma attacks, doctors can quickly provide emergency treatment to the patient.
Agriculture
The use of IoT in agriculture brings about numerous advantages. This section highlights some of the most noticeable benefits [40].
  1. 1.
    Community Farming
    The implementation of IoT can facilitate the development and advancement of community farming. Through IoT, it becomes feasible to support services that enable shared data storage, the exchange of data and information, and the enhancement of communication among farmers and agriculture experts [41]. Additionally, IoT facilities or mobile phones can be used to share various agricultural equipment within the community, either for free or through paid services.
  2. 2.
    Safety Control and Fraud Prevention
    In agriculture, the challenge extends beyond just increasing production; it also involves ensuring the production of safe and nutritious food. There have been instances of food fraud, including adulteration, counterfeiting, and artificial enhancements, which pose significant health risks and negative economic impact [42]. Researchers in [43] have explored various aspects of food fraud, including product integrity, process integrity, people integrity, and data integrity, which can be effectively addressed through IoT technology. Additionally, IoT can be integrated into both logistics traceability and the qualitative tracking of food [44].
  3. 3.
    Wealth Creation and Distributions
    The adoption of IoT will uncover novel business models, allowing individual farmers to directly connect with customers, eliminating the need for intermediaries. Consequently, this will lead to increased profits for the farmers.
  4. 4.
    Cost Reduction and Wastage
    A significant benefit of IoT is its capability for remote monitoring of devices and equipment. Farmers can oversee extensive fields from a distance, eliminating the need for physical visits and saving time. Additionally, precise knowledge of when and where to apply pesticides and fertilizers will reduce costs and minimize material wastage [45].
  5. 5.
    Operational Efficiency
    IoT technology brings operational efficiency not only to farmers but also to decision-makers in the agriculture sector, including government and non-governmental agencies. By collecting data through IoT-based agriculture surveillance, this information can guide various interventions such as disease prevention, wildfire management, compensation programs, and resource allocation.
    Farmers can leverage IoT and data analytics for making informed decisions about farm management and processes. Automated health monitoring of livestock or crops enables efficient diagnosis and prescription of medicine by veterinary or agriculture officers, reducing losses. Additionally, IoT optimizes the agri-food supply chain, ensuring real-time balance between demand and supply.
  6. 6.
    Awareness
    IoT is anticipated to create cost-effective mobile apps and wireless network services for the agriculture industry. This development can make information about product markets, prices, government services, and regulatory standards for various farm produce easily accessible. As a result, people can receive notifications when fresh products hit the market and connect with farmers interested in organic foods.
  7. 7.
    Asset Management
    IoT simplifies asset management by enabling remote monitoring of farm equipment, preventing theft, and issuing alerts for necessary maintenance.
Education
The educational sector can derive similar benefits from the utilization of IoT, much like the advantages observed in other industries such as healthcare or retail, as previously described [34]. Some common examples are presented below.
  1. 1.
    Improvement in Learning Environment
    IoT integration has the potential to revolutionize the learning experience for all students. One innovative approach involves attaching RFID tags to various objects, allowing students to interact with these objects and instantly access relevant information. This not only promotes greater student engagement but also makes learning a more enjoyable and interactive process.
  2. 2.
    Distance Learning
    IoT has the potential to enhance distance learning, enabling students facing physical classroom attendance challenges, whether due to health concerns or other reasons, to actively engage in virtual classrooms.
  3. 3.
    Enhanced Assistant for Special Students
    IoT makes learning more accessible for special students who require additional support and assistance. As referenced in Selinger et al. [46] and Lutz et al. [47], IoT is confident to offer minority and disabled learners the opportunity for high-quality education and peer-to-peer interaction. It will also facilitate student-to-teacher engagement, ultimately enhancing their prospects for greater academic success.
Warehouse
The implementation of IoT in warehouse management and processing has paved the way for smart warehousing. As a result, big industries like Alibaba, Amazon, and DHL have already integrated this technology into their inventory, logistics, and warehouse management systems. Here are descriptions of several common benefits of smart warehousing:
  1. 1.
    Efficiency and Transparency
    The primary and most essential requirement in inventory and warehouse management is to ensure efficiency and transparency across the entire process. Barreto et al. [48] have presented the key aspects of Logistics 4.0 where they propose the use of an intelligent system and a database to facilitate data sharing through an IoT platform. This interconnected approach integrates all warehouse operations and workforces, leading to increased transparency in the process and enhanced system efficiency.
  2. 2.
    Efficient Order Monitoring
    Implementation of IoT can help in monitoring the order without any human intervention. The authors of [49] has created a system that employs the following components: (1) Bluetooth tags for transmitting relevant product data; (2) a receiver for monitoring the movement of objects and the position of pallets; and (3) an integrated software package and database server for warehouse monitoring and data exchange. This developed system significantly enhances the efficiency of stock management, delivery timing, tracking, and product management.
  3. 3.
    Ensure Quick Tally
    In the traditional tally process, which relies entirely on manual labor, the task of updating product numbers and information is labor-intensive. However, through IoT integration, the tally process can be significantly expedited by simply scanning the barcodes on the products [50].
  4. 4.
    Enhance Warehouse Automation Management
    IoT technology can enable warehouse workers to inspect storage boxes without the need for physical contact, a departure from the traditional method of physically touching each box to check its contents.
Supply Chain Management
The supply chain represents a intricate logistics system comprising services that transform raw materials into finished products and subsequently deliver them to customers. Furthermore, Ensuring a continuous and on-demand supply of products is a crucial market requirement. IoT introduces groundbreaking advantages in supply chain management, some of which are detailed below [38].
  1. 1.
    Decision Making
    In the field of logistics, the frequency of information flow and the need for instant decision-making surpasses that of most other industries due to its unique nature. Sensors integrated with products facilitate the real-time and continuous generation of data, simultaneously allowing the analysis of this data to predict market trends. This predictive capability aids in making efficient decisions across the entire supply chain.
  2. 2.
    Smart Transportation
    The integration of IoT technology facilitates intelligent transportation solutions, encompassing the monitoring of delivery vehicles, tracking vehicle locations, and promoting safe driving practices. By employing sensors connected to all delivery vehicles, real-time tracking of vehicle locations becomes possible, and notifications can be dispatched in the event of accidents. Additionally, these sensors can monitor factors like excessive vehicle speed, fuel tank levels, and tire pressure to ensure safe and efficient transportation. Moreover, by analyzing traffic conditions and road conditions, the system can send notifications to the driver regarding the status of the road leading to their destination. Additionally, the system can suggest alternative routes that promote safe driving.
  3. 3.
    Smart Loading and Unloading
    Efficiently handling the loading, unloading, and stacking of products is crucial for item distribution. It is essential to achieve this with cost-efficiency and timeliness. Automation lifts equipped with RFID tags can streamline this process by precisely locating and delivering goods without requiring human intervention. Furthermore, robots can be employed as warehouse workers, leading to a reduction in human labor [51].
  4. 4.
    Smart Carriage
    Integrating Automatic Guided Vehicles (AGV) into both internal and external transportation systems can lead to cost savings and time efficiency gains [52]. Various communication technologies facilitate internal communication among these vehicles, contributing to more efficient management processes as a result [53].
  5. 5.
    Smart Packaging
    IoT has seamlessly replaced traditional packaging processes with intelligent and smart packaging solutions. These smart packages are capable of making real-time decisions, and product information is easily accessible through smart labels affixed to the items [54]. In [55], authors have employed a similar process to categorize products based on their temperature sensitivity. Furthermore, to ensure the safe delivery of products, an automatic reminder system have been proposed that activates if the products deviate from the preset conditions.
  6. 6.
    Smart Distribution
    In the logistics and warehouse management of product distribution, several stages are involved both before and during the process. Before the final goods are dispatched, there is a crucial step of labeling products according to their types and ensuring proper label attachment. This task is efficiently handled by an autonomous wheeled robot [56]. In specific industries like the food sector, smart labeling tags equipped with microcontrollers are employed to effectively track food items [57]. IoT plays a significant role in various aspects of the distribution process, including route planning, item location, sorting, and labeling.
1.4 IoT Architecture
Architecture refers to a structured framework depicting the tangible components of a network, their operational arrangement and setup, underlying principles and structures, as well as the manner in which data is organized and utilized within its functioning. IoT architecture comprises of a collection of devices, sensors, actuators, end users, cloud services and most importantly various communication layers and IoT protocols. Owing to the dynamic characteristics of the IoT domain and the crucial importance of sensors, it is impossible to identify one fundamental and universally agreed architecture for diverse IoT domains. Nevertheless, certain foundational elements will exhibit similarity across various projects.
The architecture of the IoT has been categorized based on a range of factors. Certain researchers have undertaken a division grounded in distinct layers, while others have pursued a division predicated on the specific objectives of the IoT project. However, there are some fundamental blocks present in more or less in every IoT architecture which provides scalability and flexibility in IoT architecture [58].
  1. 1.
    Sensing Layer
    The primary function of this layer involves real-time detection of changes in the connected devices. This layer comprises two main actors: sensors and actuators. Sensors are responsible for acquiring diverse data types from the environment, including but not limited to temperature, humidity, caloric information, heart rate, and object identification. Meanwhile, actuators are mechanized components tasked with device manipulation, encompassing actions such as activating and deactivating door lights and similar functionalities.
  2. 2.
    Communication Layer
    Functioning as an intermediary system bridging the sensing layer with the upper layers, this layer manages connectivity, facilitates data transmission between the sensing and cloud layers, or upper layers, and conducts message routing among remote devices.
  3. 3.
    Cloud Layer
    Situated at the core of the IoT architecture, this layer functions as the central processing unit. It receives data collected from the sensing layer to conduct storage, processing, and analysis procedures. Within this framework, the cloud employs a data centre as a pivotal server for tasks like data refinement by eliminating redundant and noisy information, data format conversion, and more. Nonetheless, the advancement of cloud computing research is progressively shifting towards the decentralisation of multiple processing tasks from the cloud to edge nodes that are currently concentrated in data centres [59].
  4. 4.
    Management Layer
    Tasked with monitoring and administering the entire IoT system, this layer handles responsibility for device configuration and firmware updates, as well as providing analytics and generating reports. These capabilities empower system administrators to comprehensively grasp the system’s dynamics and find opportunities for enhancement.
  5. 5.
    Services and Application Layer
    This layer provides value to the end user of an IoT system. This layer utilizes processed data and provide various functionalities or services.
  6. 6.
    Security Layer
    With the responsibility of ensuring security and safeguarding against unauthorized access, this layer bears the duties of authentication and authorization, intrusion detection, threat mitigation, and related responsibilities [60].
Figure 1.3: Generic architecture of IoT
The IoT architecture can be divided in two ways: (1) Layered Architecture and (2) Domain-Specific Architecture. This section presents several existing architectures for IoT. However, detailed explanations are beyond the scope of this chapter; instead, it provides an overview of a generalized architecture.
1.4.1 Generic Architecture of IoT
IoT follows a layered architecture, which refers to the structured framework used to design and organize the various components and functions of an IoT system. This architecture is composed of multiple layers, each with its specific role and responsibilities, allowing for efficient communication, data processing, and management within the IoT ecosystem. The typical layers in IoT architecture include the perception layer, network layer and application layer. These layers collaborate to ensure the smooth operation of IoT devices, data transmission, and the provision of IoT services to end users.
According to this generalized architecture, also known as the three layer architecture, the IoT system is divided into three layers [61] namely (i) application layer, (ii) network layer, and (iii) perception layer. Every one of these layers possesses inherent security challenges [62]. A visual representation of the Three Layer Architecture is featured in Fig. 1.3. The details are described below.
  1. 1.
    Perception Layer
    The perception layer, also known as the sensor layer is the foundational layer of IoT architecture [61]. This layer engages with smart devices, including but not limited to smartwatches and smart rings, employing an array of sensors and actuators. The principal objective of this layer pertains to the collection of data from these intelligent devices via sensors, subsequently transmitting the acquired data to the upper layer known as the network layer.
  2. 2.
    Network Layer
    The network layer, also known as the transmission layer is the middle layer of IoT architecture [63]. This layer is responsible for receiving the information passed from the perception layer and determine the routes to transmit the processed data to various connected IoT devices and applications using integrated network such as wired or wireless secure connection. Network layer is core layer of IoT three layer architecture as this uses various devices such as routing devices, gateways, switches and hub and operate them by using various communication technologies such as WiFi, Bluetooth, 3G, LTE, Zigbee etc. In summary, the network layer is responsible for transmitting data to and from several applications through interfaces and gateways using multiple communication technologies and protocols.
  3. 3.
    Application Layer
    Serving as the uppermost tier within the IoT architecture, denoted as the application layer or business layer, as referenced in [58], this layer is tasked with the aggregation of data from the network layer, thereby striving to attain the objective of establishing an smart environment, the ultimate aim of the IoT paradigm. This layer accommodates a diverse array of applications, each characterised by their own requirements. Examples of such applications are smart grids, smart cities, and smart transportation, as elaborated in [62]. Moreover, this layer assumes the responsibility of upholding the data’s authenticity, integrity, and confidentiality, as elaborated in [64].

### 12 Jan 2024 v2

@@ 1.2.2 Components
1.2.3 Characteristics
This segment essentially serves as the user interface (UI). The UI offers a platform for users to interact directly with the application or system, facilitating communication. A user interface is not always reliant on a screen. For instance, a TV remote utilizes a user interface comprising multiple buttons, while devices like the Amazon Echo respond to voice commands for control. Receiving automatic notifications, monitoring information proactively, and controlling the system remotely are some common examples of user interfaces in IoT systems [30]. The IoT characteristics discussed in this chapter are illustrated in Figure 1.1.
@@ 1.2.3 Characteristics
    Connectivity: Connectivity is the most vital requirement of IoT. The main aspect of IoT is a network with millions of devices connected to each other. The connectivity remains constant, allowing anyone from anywhere to connect to the IoT network at any given moment.
@@
    Intelligent decision making capability: The extraction of knowledge from the data generated is highly significant. Consider a sensor that produces data; however, the true value of that data lies in its proper interpretation. This represents a crucial aspect of IoT, wherein IoT devices possess the capability to transform raw data collected by sensors into meaningful information and make decisions based on it.
@@
    Dynamic and Self Adapting: IoT devices should have the ability to adapt to changes in context, their surrounding environment, and the existing situation. For instance, within a surveillance system, cameras are capable of switching between day and night modes or adjusting their resolution in response to motion detection, demonstrating their adaptability.
@@
    Unique Identity: Every IoT device should have a unique identity and unique identifier. IoT device interfaces enable users to inquire about device information, monitor their status, and remotely manage them. Having a distinct identity is essential to empower users to safeguard their devices, whether through password protection or alternative security measures [33].
@@
    Self Configuring: IoT devices possess the capability to autonomously update their systems in response to the situation, eliminating the need for user intervention. Moreover, they exhibit flexibility in network management, allowing new devices to seamlessly join the network and permitting any device to depart from the network at any time.
@@
    Scalability: The IoT network is experiencing a continual growth in the number of connected devices, resulting in a substantial and continuous generation of data. Consequently, scalability emerges as the foremost feature of any IoT system.
    Figure 1.2: Advantages of IoT
@@
    Interoperability: IoT devices rely on standardized protocols and technologies to guarantee seamless communication among themselves and with other systems. Interoperability represents a fundamental block of the IoT, signifying the capacity for various IoT devices and systems to interact and share data, irrespective of the underlying technology or the manufacturer. Hence, IoT devices employ standardized protocols, data formats, and technologies to uphold interoperability.
@@
    Heterogeneity: The devices within an IoT network exhibit heterogeneity, showcasing the network’s capability to accommodate diverse elements.
@@
    Energy Efficiency: Energy efficiency stands out as a significant characteristic of IoT. Numerous IoT devices are purposefully crafted to minimize energy consumption and create low-power devices. Furthermore, various IoT approaches are tailored to optimize power usage, such as opting for fog/edge computing over cloud computing to decrease bandwidth requirements and reduce power consumption when constructing an IoT system [34].
@@
    Security: The presence of millions of internet-connected devices and the vast amount of data generated underscore the vulnerability of the IoT network to security threats. Hence, safety and security emerge as pivotal characteristics of the IoT. Ensuring safety is of paramount importance to preserve the effectiveness of IoT’s advantages, including efficiency and novel experiences.
@@ 1.3 Advantages of IoT
The amount of benefits IoT offers to people’s lives is the main reason IoT is becoming more popular by the day. IoT solutions are developed and invented to make people’s lives easier and more convenient. IoT technology affects almost every field, including healthcare, education, and business. This section elaborates on the most visible benefits of IoT in people’s daily lives [35, 36], and Fig. 1.2 provides a visual overview.
@@
    Efficient Data Collection: IoT-based data collection has been particularly beneficial in sectors such as healthcare and finance [37]. A straightforward illustration of efficient data collection is the integration of IoT in the retail sector. Internet-connected tags can provide data on purchase decisions and sales trends, whether weekly or monthly. This enhanced data gathering can improve inventory management and reveal valuable insights into customer behavior, ultimately contributing to the prosperity of the business.
@@
    Control and Automation: IoT has provided its customers with a more convenient lifestyle and allows them to control their daily activities with the touch of a button. One simple yet remarkable example is the smart bulb, which can be controlled without even touching the switches; the user can simply turn off or on the light remotely. Not only can the light bulb, coffee maker, or any other electrical device in the home be controlled with the tip of the device, but they can also be controlled using a voice command when they are linked to Google’s or Amazon’s voice assistants.
@@
    Real-time Access to Information: IoT devices offer immediate access to information, proving invaluable across healthcare, business, and everyday applications. A prime illustration of the advantages of real-time data access is within the healthcare sector. Physicians can access patient data in real-time, enabling continuous health monitoring. This capability becomes particularly crucial in delivering emergency medical assistance swiftly when unexpected health issues arise.
@@
    Improved Efficiency: IoT systems operate autonomously, which is a valuable asset in various domains. Reduced human intervention leads to increased efficiency and decreased labor reliance. For instance, a company with a fleet of delivery vehicles can effortlessly track their real-time locations, eliminating the need for manual employee involvement in this task.
@@
    Improved Quality of Life: The advent of IoT has significantly improved the lives of its users in numerous ways. Real-time health monitoring, including devices like blood pressure monitors and fitness trackers, empowers users to maintain their well-being effectively. Smart homes offer a stress-free and effortless lifestyle. These advantages extend beyond individuals and can benefit entire industries or communities. Smart devices, linked not only to intelligent traffic lights but also to road safety monitors and toll gates, can provide drivers with real-time information about road conditions on their route.
@@
    Cost and Time Effective: IoT minimizes human effort and relies heavily on real-time data transmission, leading to time savings. For example, real-time patient monitoring benefits both patients and doctors by eliminating the need for physical meetings, thus saving time for both parties. IoT aids businesses in streamlining their workflows by offering valuable insights and real-time information, resulting in cost reductions. In addition to businesses, individuals can reduce their everyday expenses through the use of IoT.
@@
    Asset Tracking: This process involves tracking products within a business or logistics management system. Manual asset tracking is labor-intensive and time-consuming, but it can be streamlined through the application of IoT technologies like barcodes and RFID tags. These technologies allow for remote monitoring of goods and provide stakeholders with information about any faults or problems in real time [38].
@@
    Data Analysis for Research Purpose: The enormous amount of data collected from IoT devices has become a blessing for researchers in various fields like healthcare, education, business, etc. The healthcare researchers can use the data collected via biosensors to invent cures and vaccines for disease; the finance industry can use the data to understand trends and improve customer experience; the super shops can analyze customer behavior and improve their businesses; and so on.
@@
    Big Data and Predictive Analysis: “Big data” has been a widely recognized term in the world long before the emergence of IoT. It involves the collection and analysis of massive volumes of data. One of the primary objectives of IoT is to amass data from diverse sources, sending this information back to systems for analysis. Effective analysis of big data can yield valuable insights, spanning from stock market predictions to understanding customer behavior, thereby enhancing the business landscape.
@@
    Increased Productivity: IoT utilization in both industry and homes has the potential to boost productivity significantly. For instance, in a smart home, users can streamline various household tasks using voice commands, enabling efficient multitasking. Similarly, in a business setting, analyzing customer behavior can enhance customer satisfaction, ultimately contributing to the prosperity of the enterprise. As an example, 46% of businesses that embraced IoT strategies saw improvements in efficiency, even though only 29% initially anticipated such improvements [39]. In the healthcare sector, doctors can offer more extensive services to their patients if they do not need to make physical visits to each patient individually.
@@
    Safety and Security: The incorporation of IoT offers users a means of security not only in their homes but also in their businesses, schools, offices, and virtually anywhere. Individuals can remotely monitor their valuable assets, such as vehicles and other important items. Parents can even keep track of their children’s whereabouts from their workplaces, providing peace of mind. IoT allows for vehicle tracking and the setup of alert systems in case of unusual incidents. Financial companies and banks can enhance the security of their confidential rooms or vehicles by utilizing IoT [34].
@@
    Enhance Customer Engagement: The IoT offers several avenues for enhancing customer engagement. It achieves this by leveraging valuable customer data, personalizing experiences, improving convenience, and enabling real-time interactions.
@@
    Efficient Resource Utilization:IoT facilitates efficient resource utilization through various mechanisms and capabilities. Its ability to collect and monitor real-time information allows organizations to track the status of their resources, such as equipment and machinery, enabling the identification of inefficiencies. Predictive maintenance powered by IoT in the industrial sector can anticipate machinery failures, reducing downtime and optimizing resource allocation for maintenance.
@@
    Improved Technology: The innovation brought about by IoT leads to the creation of newer and more advanced technologies on the market. For instance, consider the scenario of an air conditioner that was initially controlled manually with a remote. With the advent of IoT, users can now operate it using voice commands or control it remotely. When such innovations hit the market, they spark competition, driving the development of improved solutions based on user feedback. This cycle of innovation in response to IoT advancements contributes to the continuous progression of technology [37].
1.4 IoT Architecture
Architecture refers to a structured framework depicting the tangible components of a network, their operational arrangement and setup, underlying principles and structures, as well as the manner in which data is organized and utilized within its functioning. IoT architecture comprises a collection of devices, sensors, actuators, end users, cloud services, and most importantly, various communication layers and IoT protocols. All IoT systems inherently follow the generic three-layer architecture. However, based on necessities or specific application requirements, the generic model can be modified by adding extra layers, thus forming four- or five-layer architectures [4].
Figure 1.3: Generic architecture of IoT
The IoT architecture can be divided in two ways: (1) layered architecture and (2) domain-specific architecture. In this section, we have presented a generic architecture of IoT. IoT follows a layered architecture, which refers to the structured framework used to design and organize the various components and functions of an IoT system. This architecture is composed of multiple layers, each with its own specific role and responsibilities, allowing for efficient communication, data processing, and management within the IoT ecosystem. The typical layers in the IoT architecture include the perception layer, network layer, and application layer. These layers collaborate to ensure the smooth operation of IoT devices, data transmission, and the provision of IoT services to end users.
According to this generalized architecture, also known as the three-layer architecture, the IoT system is divided into three layers [61], namely (i) the application layer, (ii) the network layer, and (iii) the perception layer. Every one of these layers possesses inherent security challenges [62]. A visual representation of the generic three-layer architecture is featured in Fig. 1.3. The details are described below.
@@ 1.4 IoT Architecture
    Perception Layer: The perception layer, also known as the sensor layer, is the foundational layer of IoT architecture [61]. This layer engages with smart devices, including but not limited to smartwatches and smart rings, employing an array of sensors and actuators. The principal objective of this layer pertains to the collection of data from these intelligent devices via sensors, subsequently transmitting the acquired data to the upper layer known as the network layer.
@@
    Network Layer: The network layer, also known as the transmission layer, is the middle layer of the IoT architecture [63]. This layer is responsible for receiving the information passed from the perception layer and determining the routes to transmit the processed data to various connected IoT devices and applications using integrated networks such as wired or wireless secure connections. The network layer is the core layer of the IoT three-layer architecture, as it uses various devices such as routing devices, gateways, switches, and hubs and operates them by using various communication technologies such as WiFi, Bluetooth, 3G, LTE, Zigbee, etc. In summary, the network layer is responsible for transmitting data to and from several applications through interfaces and gateways using multiple communication technologies and protocols.
@@
    Application Layer: Serving as the uppermost tier within the IoT architecture, denoted as the application layer or business layer, as referenced in [58], this layer is tasked with the aggregation of data from the network layer, thereby striving to attain the objective of establishing a smart environment, the ultimate aim of the IoT paradigm. This layer accommodates a diverse array of applications, each characterized by their own requirements. Examples of such applications are smart grids, smart cities, and smart transportation, as elaborated in [62]. Moreover, this layer assumes the responsibility of upholding the data’s authenticity, integrity, and confidentiality, as elaborated in [64].